1. Let A and B be two non-empty finite sets. If cardinalities of the sets A, B, and A∩B are respectively 72, 28 and 13, then find the cardinality of the set A∪B.

2. If n(A-B)=45, n(AUB )=110 and n(A∩B)=15, then find n(B).

3. If n(A)=33, n(B)=36 and n(C)=28, find n(A∪B∪C).
1
Expert's answer
2020-05-18T19:58:36-0400
  1. 72+28-13=87
  2. n(B-A)=110-45-15=50. Then n(B)=15+50=65
  3. If these sets are disjoint, then 33+36+28=97